A realtor wants to compare the average sales-to-appraisal ratios of residential properties sold in four neighborhoods (A, B, C, and D). Four properties are randomly selected from each neighborhood and the ratios recorded for each, as shown below.

A:        1.2, 1.1, 0.9, 0.4                C:   1.0, 1.5, 1.1, 1.3

B:        2.5, 2.1, 1.9, 1.6                D:   0.8, 1.3, 1.1, 0.7

Interpret the results of the analysis summarized in the following table:

Source                 df        SS             MS            F        PR > F

Neighborhoods             3.1819     1.0606       10.76      0.001                                              

Error                    12           

Total                             4.3644

a) What is the among group degrees of freedom?

12

16

3

4

IS 12 CORRECT
Answer #1

A realtor wants to compare the average sales-to-appraisal ratios of residential properties sold in four neighborhoods (A, B, C, and D).

There are 4 treatments.

k = 4

a) The among group degrees of freedom = k - 1 = 4 - 1 = 3